40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon May 18 19:44:52 2009, rana, steve, Configuration, VAC, Cryo Pump turned off and valved off: 1 beer can only 
    Reply  Mon May 18 20:16:20 2009, rana, Configuration, VAC, not it a.png
    Reply  Tue May 19 09:34:29 2009, steve, Configuration, VAC, IFO is not pumped & CRYO is being regenerated cryoreg.jpgcryo30d.jpg
       Reply  Tue May 19 16:18:45 2009, steve, Configuration, VAC, IFO pressure is 4.2e-7 Torr with CRYO pump 
Message ID: 1602     Entry time: Mon May 18 20:16:20 2009     In reply to: 1601
Author: rana 
Type: Configuration 
Category: VAC 
Subject: not it 
There was essentially no change in the ETMY oplev spectrum with the cryo off!

So I went out to the ETMY OL table to see what else was going on. I found there one of
the most horrible opto-mechanical setups
I have ever seen (and remember that I have once
seen someone mount an NPRO on a cardboard box). Some bad person had mounted the ETMY OL
lens on a 12" long skinny post and extended it towards the viewport. So there was a post
holder on the table and a lens ~12" away on a rickety lever arm.

I took this lens away and the spectrum is now good. Shame on you.
CYAN -  cryo ON
BLACK - cryo OFF
BLUE -  no crappy lens + mount

This OL needs to be fixed correctly by putting in a proper lens to get a small spot on the QPD.
Attachment 1: a.png  83 kB  | Hide | Hide all
a.png
ELOG V3.1.3-